# What does a regular star look like?

### v1

@@ I Introduction
Over the past decade, many predictions of general relativity have been confirmed in astronomical observations and laboratory experiments. In 2015, LIGO and Virgo collaborations detected gravitational waves from merging black holes, directly confirming both phenomena [1, 2, 3, 4, 5, 6, 7]. In 2019, the Event Horizon Telescope (EHT) released the first image of a black hole in the galaxy M87 and later imaged the black hole at the center of our own galaxy, the Milky Way. These achievements allowed for direct observation of black holes with accretion disks [Imagin1, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Research on gravitational waves and black holes earned Nobel Prizes in Physics in 2017 and 2020, respectively.
@@ II Null Geodesics in regular star spacetime
For photons with $r_{m}=r_{n}$, the trajectory also includes the outgoing portion after passing the periastron, which is symmetric with respect to the incoming portion about the periastron. According to the definition of the number of orbits $n\equiv\phi/2\pi$ in Gralla (2019) [8] and the above equation, the corresponding formula for $n$ is given by Eq.(10). The second row of Fig.1 shows the relation between $n$ and $b$ for different values of $\beta$.
@@ III.2 Geometrically Thin and Optically Thick Disk Emission
We also present the images obtained under the geometrically thin and optically thick disk model. The ray-tracing treatment of this accretion-disk model follows that of standard black hole accretion-disk modelsLABEL:luminet_1979,tian_2019,zhang_2021,hou_2022,meng_2023,guo_2024. The coordinate system is constructed as illustrated in Fig.5, where the deflection angle $\gamma$ obeys the following mathematical relationship with the observer’s inclination angle $\theta_{0}$ and $\alpha$ [36, 37]
@@ IV Conclusions and discussion
In this work, we investigate the images of accretion disks around regular stars. We find that, in such spacetimes, photons can pass through the stellar center, giving rise to a central bright spot. This result is consistent with recent astronomical observations [26]. We study representative accretion disk models in this spacetime and find that, in addition to the transparency of the stellar center, the secondary images of the accretion disk become incomplete when the observation angle is sufficiently large. This feature may serve as a decisive observational signature for future astronomical observations aimed at identifying such regular stars.

### v2

@@ I Introduction
Over the past decade, many predictions of general relativity have been confirmed in astronomical observations and laboratory experiments. In 2015, LIGO and Virgo collaborations detected gravitational waves from merging black holes, directly confirming both phenomena [1, 2, 3, 4, 5, 6, 7]. The Event Horizon Telescope (EHT) released the first image of a black hole in the galaxy M87 [8, 9, 10, 11, 12, 13] in 2019 and later imaged the black hole Sagittarius A* [14, 15, 16, 17, 18, 19] at the center of our own galaxy, the Milky Way in 2022. These achievements allowed for direct observation of black holes with accretion disks [20]. Research on gravitational waves and black holes earned Nobel Prizes in Physics in 2017 and 2020, respectively.
@@ II Null Geodesics in regular star spacetime
For photons with $r_{m}=r_{n}$, the trajectory also includes the outgoing portion after passing the periastron, which is symmetric with respect to the incoming portion about the periastron. According to the definition of the number of orbits $n\equiv\phi/2\pi$ in Reference [20] and the above equation, the corresponding formula for $n$ is given by Eq.(10). The second row of Fig.1 shows the relation between $n$ and $b$ for different values of $\beta$.
@@ III.2 Geometrically Thin and Optically Thick Disk Emission
We also present the images obtained under the geometrically thin and optically thick disk model. The ray-tracing treatment of this accretion-disk model follows that of standard black hole accretion-disk models [35, 36, 37, 38, 39, 40]. The coordinate system is constructed as illustrated in Fig.5, where the deflection angle $\gamma$ obeys the following mathematical relationship with the observer’s inclination angle $\theta_{0}$ and $\alpha$ [35, 36]
@@ IV Conclusions and discussion
The study of non-singular black holes and black hole shadow is of current interest [42, 43, 44, 45, 46, 47, 48, 49, 50] . In this work, we investigate the images of accretion disks around regular stars. We find that, in such spacetimes, photons can pass through the stellar center, giving rise to a central bright spot. This result is consistent with recent astronomical observations [27]. We study representative accretion disk models in this spacetime and find that, in addition to the transparency of the stellar center, the secondary images of the accretion disk become incomplete when the observation angle is sufficiently large. This feature may serve as a decisive observational signature for future astronomical observations aimed at identifying such regular stars.